OpenStudy (anonymous):

The ratio of Bills money to Henry's money was 5:6 After Bill spent $800 on a TV set, the ratio became 1:2. How much money did Bill have at first?

OpenStudy (anonymous):

b/h = 5/6 is 6b = 5h (b-800)/h = 1/2

OpenStudy (anonymous):

6b = 5h h = 2b - 1600 6b = 5(2b - 1600) b = 2000

OpenStudy (anonymous):

set both equations up as ratios and then cross multiply using substitution you can find B

OpenStudy (anonymous):